Update for scheduling physicist -- A late Happy Halloween and what's going on this weekend.

From: Flemming Videbaek (videbaek@sgs1.hirg.bnl.goV)
Date: Fri Nov 02 2001 - 20:28:50 EST

  • Next message: Apache: "Shift report 20011102 16:00-24:00"

    FYI
    
    I sincerely hope the coll at brahms gets straightened out - yesterday was just awfull for us.
     (great for rhic and phenix though in reaching luminosity goals)
    
    Flemming
    
    ------------------------------------------------------
    Flemming Videbaek
    Physics Department
    Brookhaven National Laboratory
    
    tlf: 631-344-4106
    fax 631-344-1334
    e-mail: videbaek@bnl.gov
    ----- Original Message ----- 
    From: Angelika Drees 
    Sent: Friday, November 02, 2001 5:09 PM
    Subject: A late Happy Halloween and what's going on this weekend. 
    
    
    Hi, 
    another friday ... First a brief summary of yesterdays machine development: it took (just a little bit) more than the promised  3 ramps but we finally made it to beta*=1m at PHENIX :). Late evening yesterday we had the first store with this lattice and got a luminosity of approximately 4*10^26 cm^-2 s^-1 at the beginning of the store (in PHENIX) and it scaled with a factor of 2 compared to the other experiments as expected. However, we sacrificed BRAHMS :( at this point and steering BRAHMS has the highest priority for the next store(s). Currently RHIC is at injection and we are preparing a ramp. 
    
    The weekend (and this evening/night) should be spent on production fills with only minor tuning in the land of lifetime improvements and local decoupling (parasitic). Monday  morning, pending a RHIC store, we will dedicate ~2 hours for injection studies with a beta*=3m lattice (preparation for the proton run), a left-over from the development day yesterday. 
    
    SEB will run over the weekend as usual with the following change in priorities: D-line  has the highest priority and should get its maximum of 25 Tp, C-line whatever remains while maintaining 6 Tp on the C' target.  We will run like this until Nov. 8th. There was a temperature problem with the D-line target earlier causing it to trip on intensities around 22 Tp. The threshold was raised today but the intensity should be lowered accordingly in case of ongoing problems.  Later tonight, when we have RHIC filled, a flag in the C-line needs to be pulled out manually (hopefully *short* AGS access required) before we can deliver SEB. 
    
    I hope you all got lots of candy on Halloween ;). Have a nice weekend, 
    
                                                    \|/
                     Angelika                      (.~.)
                                                  o0(_)0o
    -----------------------------------------------------------
    Angelika Drees (Collider-Accelerator Department, AP group)
    BNL, Building 911b-210
    work: (631) 344 2348
    FAX:  (631) 344 5954      
    home: (631) 474 3884
      
    



    This archive was generated by hypermail 2b30 : Fri Nov 02 2001 - 20:24:51 EST